# Masked Image Modeling via Dynamic Token Morphing

### v1

@@ Abstract
Masked Image Modeling (MIM) arises as a promising option for Vision Transformers among various self-supervised learning (SSL) methods. The essence of MIM lies in token-wise masked patch predictions, with targets patchified from images; or generated by pre-trained tokenizers or models. We argue targets from the pre-trained models usually exhibit spatial inconsistency, which makes it excessively challenging for the model to follow to learn more discriminative representations. To mitigate the issue, we introduce a novel self-supervision signal based on Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens. DTM can be generally applied to various SSL frameworks, yet we propose a simple MIM that employs DTM to effectively improve the performance barely introducing extra training costs. Our experiments on ImageNet-1K and ADE20K evidently demonstrate the superiority of our methods. Furthermore, the comparative evaluation of iNaturalist and Fine-grained Visual Classification datasets further validates the transferability of our method on various downstream tasks. Our code will be released publicly.
@@ 1 Introduction
Since the success of Vision Transformer (ViT), numerous training strategies have arisen for ViT [13], including self-supervised learning (SSL) methods [8, 16, 15, 6]. Recent advances in masked image modeling (MIM) [47, 17] have established it as a mainstream SSL approach for ViT. In essence, the crux of the MIM methods lies in leveraging token-wise optimization objectives derived from masked patch prediction. Consequently, MIM methods have explored various approaches to generate effective target patches, employing various pre-trained models including vision-language models [4, 28], utilizing momentum teachers [3, 51] or directly exploiting patchified images [17, 47].
@@
However, we observe that patch representations generated by the pre-trained models tend to show spatial inconsistency (See §3). Specifically, for example, the patch-wise zero-shot classification of a pre-trained vision-language model reveals spatially inconsistent prediction results, as shown in Fig. 2. To delve deeper into the behaviors of patch representations from pre-trained models, we analyze the impacts of spatial inconsistency from the perspective of discriminability and training capability. We first observe that patch-wise ensemble prediction reveals inferior discriminability in zero-shot image classification and linear probing without token morphing, which aggregates contextually similar tokens. Furthermore, our distillation experiments reveals that token morphing without contextual awareness disrupts pre-training, whereas the model pre-trained using token morphing with context enjoys the potential to enhance representation quality. Consequently, spatially inconsistent target patches pose challenges in learning representations.
@@
Motivated by the observations, we introduce a novel masked image modeling method with our Dynamic Token Morphing (DTM). We conjecture that training should be accelerated by the guidance of composite representations derived from morphed tokens, which aggregate contextually related tokens (see Fig. 1). More specifically, we encode the token-wise target representations and derive matching relations among tokens through DTM. The merging process is performed for both sides of tokens regarding the matching relation and aligning each morphed token with the corresponding merged target tokens. The range of our morphed tokens can vary from a single token to the whole tokens, covering from token-wise to image-level representation learning. We employ the bipartite matching algorithm for token morphing in the aspects of efficiency and efficacy.
@@
Through extensive experiments, we verify the general applicability and scalability of our DTM. We observed that our method improves fine-tuning accuracies across various types of SSL frameworks, affirming its general applicability on SSL methods. Moreover, DTM enhances the baseline from ViT-S/16 to ViT-L/16, demonstrating its scalability. Comparison with previous SSL methods verifying its superiority. Our method shows improved transferability on the iNaturalist and Fine-grained Visual Classification datasets.
@@ 2 Related Work
Masked Image Modeling.
@@
Inspired by the promising performance of masked language modeling, BEiT [4] successfully extends the MIM framework into the computer vision domain, using an external offline tokenizer from Dall-E [30]. iBOT [51] jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec [3] incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from multiple layers of neural networks. MAE [17] and SimMIM [47] demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, there have been several attempts to exploit the pre-trained model as a tokenizer. BEiT v2 [28] pre-trains a codebook for CLIP [29] to discretize a semantic space. MVP [43] exploits a tokenizer pre-trained with multimodal data to enhance semantics for MIM. A line of studies [44, 33] using CLIP as a teacher to use it as target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
Token Aggregation Methods.
@@
Token aggregation has mainly been employed to enhance efficiency. Clustering is one of the commonly used token aggregation approaches. Hard clustering methods like K-Means [24], K-Medoids [1], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) enforce each data to exclusively belong to a single cluster. Bipartite matching [19] also aggregates data in a hard clustering manner, optimizing pairs of elements from two disjoint sets given objective function. Soft clustering allows data to belong to multiple clusters. LIT [27] employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods [31, 48, 35, 23] can be categorized as token aggregation methods; however, they have deeply focused on compressing tokens to achieve a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We leverage the concept of token aggregation to address spatial inconsistency in patch-level supervision, thereby enhancing the effectiveness of masked image modeling in terms of precision.
@@
| Token morphing | Zero-shot image classification () | Linear probing () |
| 26.5 | 73.2 | |
| ✓ | 30.8 (+4.3) | 77.6 (+3.2) |
3 Motivation
In this section, we study the spatially inconsistent behavior of patch representations generated by pre-trained models and its impacts on pre-training. We visualize the spatial inconsistency of patch representations from a vision-language model in Fig. 2. Then, we conduct a quantitative analysis of the spatial inconsistency on the dataset scale in Table 1. Finally, we compare the distillation capability among token-wise, down-sampling-based, and several token aggregation-based objectives to confirm the efficacy of representation learning via context-aware token morphing.
Visualization of spatial inconsistent predictions.
To verify the spatial inconsistency among patch representations of the pre-trained model, we attempt to investigate spatially inconsistent predictions among patches through patch-wise zero-shot classification. To emphasize the spatial inconsistent representations, we conduct a comparative analysis between vanilla patch-wise prediction and prediction with smoothed tokens through token morphing, which aggregates contextually related tokens. Given the input image in Fig. 1(a), we visualize the patch-wise prediction results with and without token morphing in Fig. 1(b) and 1(c), respectively. Here, patches with accurate predictions are colored in green, while misclassified patches are colored in red. The wrong patches gradually transition to a darker red as the predicted class deviates from the ground-truth class.
| Distillation signals | Linear probing () |
| Token-wise (baseline) | 70.9 |
| Downsampling | 68.8 (-2.1) |
| Layer-wise Bipartite Matching [5] | 70.2 (-0.7) |
| Dynamic Token Morphing (Ours) | 72.2 (+1.3) |
@@ 3 Motivation
Utilizing token morphing for predictions yields improved accuracy with 113 correct patches out of 196 and a gap of 31 patches compared to vanilla patch-wise prediction, demonstrating the importance of spatial consistency. On the other hand, despite the proximity and contextual similarity among patches, the wrong patches in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (french bulldog, gorilla, bull mastiff, hotdog, and newfoundland dog) while patches in Fig 1(c) show correct or relatively consistent predictions (french bulldog). The comparison between the two prediction results highlights the spatially inconsistent representations among patches of the pre-trained models, potentially disrupting representation learning when used as targets.
Analysis of the impact of spatial inconsistency.
To quantitatively assess the impacts of spatial inconsistency on the dataset scale, we conduct a quantitative comparison between patch-wise ensemble predictions with and without token morphing. Specifically, in the case of patch-wise ensemble prediction without token morphing, we predict the class scores for each patch and average the scores across the patches within the given image. In contrast, in the case of token morphing, we aggregate groups of semantically associated tokens and average the token representations group-wise before patch-wise ensemble prediction. We use the CLIP Base model for the encoder. We aggregate 98 tokens for token morphing, which occupy half of all tokens. Table 1 reports zero-shot image classification and linear probing results on ImageNet-1k. Patch-wise ensemble prediction with token morphing exceeds the performances without token morphing, which reveals that patch representations exhibit spatial inconsistency, and the token morphing for contextually related tokens can remedy the inconsistency.
Comparison on distillation capability varying learning objectives.
We further conduct a quantitative comparison of distillation approaches with varying configurations, as described in Fig. 3. The goal is to demonstrate the importance of resolving spatial inconsistency for representation learning while showing that indiscriminate aggregation may risk degrading the representations. We compare token-wise distillation, distillation with downsampled supervisions, distillation with layer-wise Bipartite Matching during encoding for targets, and distillation with our Dynamic Token Morphing. For each configuration, we pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K. As shown in Table 2, distillation with token morphing exceeds the performance of the token-wise distillation approach, revealing that patch representations exhibit spatial inconsistency, and token morphing for contextually related tokens can remedy this inconsistency. However, since the downsampling approach aggregates tokens without context awareness, and the layer-wise Bipartite Matching approach has a risk of harming final representations during encoding, both approaches exhibit degraded performances. The results emphasize the importance of context-aware token morphing.
4 Method
We introduce a novel masked image model training based on our Dynamic Token Morphing (DTM). In DTM, we analyze the relation between token representations and aggregate similar tokens to build morphed tokens representing various semantic information. By aligning morphed tokens from online and target encoders, the masked image model is applied in morphed token-level rather than token-level loss of conventional methods. DTM is designed for universal application across various online and target encoders. The framework of DTM is described in Fig. 4, where the DTM module is straightforwardly added to the baseline MIM scheme.
4.1 Token Encoding Process Formulation
@@
We begin by formulating the token encoding process. Given an image , we patchify the image into N patches . We select positions of masked patches in a block-wise manner with a masking ratio so that . We mask the image patches by replacing the image patches of the position in to a shared learnable embedding . To this end, the input of the encoders becomes where for and for . The masked patches are concatenated with a learnable CLS token and fed into the online encoder with a subsequent linear head and the target encoder , and become encoded online tokens and where and , respectively.
| Framework | Method | FT. () |
| MAE + CLIP | 82.6 | |
| Feature MIM | MAE + CLIP + DTM | 83.2 (+0.6) |
| (CLIP teacher) | BEiT v2 | 84.2 |
| BEiT v2 + DTM | 84.4 (+0.2) | |
| Image-level SSL | BYOL | 81.7 |
| BYOL + DTM | 82.1 (+0.4) |
4.2 Dynamic Token Morphing
Considers -th DTM among total levels DTM, dynamic token morphing scheduling randomly assigns number of final morphed tokens and number of iterations to aggregate tokens. Then, we compute token morphing schedule , a sequence of token numbers to aggregate. Rather than samples a sequence of random numbers that satisfy , we simply define as:

### v2

# Morphing Tokens Draw Strong Masked Image Models

@@ Abstract
Masked image modeling (MIM) is a promising option for training Vision Transformers among various self-supervised learning (SSL) methods. The essence of MIM lies in token-wise masked token predictions, with targets tokenized from images or generated by pre-trained models such as vision-language models. While tokenizers or pre-trained models are plausible MIM targets, they often offer spatially inconsistent targets even for neighboring tokens, complicating models to learn unified discriminative representations. Our pilot study confirms that addressing spatial inconsistencies has the potential to enhance representation quality. Motivated by the findings, we introduce a novel self-supervision signal called Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens to yield contextualized targets. DTM is compatible with various SSL frameworks; we showcase an improved MIM by employing DTM, barely introducing extra training costs. Our experiments on ImageNet-1K and ADE20K demonstrate the superiority of our methods compared with state-of-the-art, complex MIM methods. Furthermore, the comparative evaluation of the iNaturalists and fine-grained visual classification datasets further validates the transferability of our method on various downstream tasks. Code is available at https://github.com/naver-ai/dtm.
@@ 1 Introduction
Since the success of Vision Transformer (ViT), numerous training strategies have arisen for ViT [15], including self-supervised learning (SSL) methods [10, 20, 18, 7]. Recent advances in masked image modeling (MIM) [52, 19, 4] have established it as a mainstream SSL approach for ViT. In essence, the crux of the MIM methods lies in leveraging token-wise optimization objectives derived from predicting masked tokens to given targets. Consequently, MIM methods have explored various approaches to generate effective target tokens, employing various pre-trained models including vision-language models [5, 34], utilizing momentum teachers [4, 56] or directly exploiting patchified images [19, 52].
@@
While tokenizers or pre-trained models have proven their potential as promising MIM targets [5, 34, 27, 49], we argue the pre-trained models tend to generate spatially noisy token representations of class information (e.g., inconsistent class labels), and their utilization as pre-training targets may incur negative impacts (See §3). For example, a pre-trained vision-language model reveals spatially inconsistent prediction results, as shown in Fig. 1. To delve deeper into the behaviors of token representations from pre-trained models, we analyze the impacts of spatially inconsistent representations from the perspective of discriminability and training capability. We initially observe reduced discriminability in zero-shot image classification and linear probing without token aggregation, also revealed with our continuous metric. Furthermore, another pilot distillation experiment reveals that the targets derived from a token aggregation method, not aware of context, disrupts pre-training. Unlikely, the model supervised by aggregated target tokens with context enjoys the potential of enhanced representation. Consequently, spatially inconsistent targets challenge learning one-to-one token maps, leading to suboptimal representation learning.
@@
Bearing those in mind, we introduce a novel token contextualization method Dynamic Token Morphing (DTM), where token morphing is a process that links contextually similar tokens and aggregates them for coherent and comprehensive representations. We conjecture that training should be accelerated by the guidance provided by composite representations of morphed tokens, derived from the aggregation of contextually related tokens (see Fig. 3). Specifically, we encode the token-wise target representations and derive matching relations among tokens through DTM. The merging process is performed for both tokens regarding the matching relation and aligning each morphed token with the corresponding morphed target tokens. The range of morphing can vary from a single token to the entire tokens, covering from token-wise to image-level representation learning. While there are many options, we simply opt for bipartite matching for morphing, achieving both efficiency and efficacy.
@@
Through extensive experiments, we verify our method’s general applicability and scalability. DTM improves fine-tuning accuracies as equipped with various SSL frameworks, affirming its general applicability as a basic element. Moreover, DTM consistently enhances the baselines from ViT-S/16 to ViT-L/16 with large margins, demonstrating its scalability. Comparison with previous SSL methods verifying its superiority. Our method showcases further transferability on the iNaturalist and fine-grained visual classification datasets.
@@ 2 Related Work
2.0.1 Masked image modeling.
@@
Inspired by the promising performance of masked language modeling (MLM), BEiT [5] successfully extends MLM into the computer vision domain, using an external offline tokenizer from Dall-E [36]. iBOT [56] jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec [4] incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from the multiple layers of neural networks. MAE [19] and SimMIM [52] demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, several attempts have been made to exploit the pre-trained model as a tokenizer. BEiT v2 [34] pre-train a codebook for CLIP [35] to discretize a semantic space. MVP [48] exploits a tokenizer pre-trained with multimodal data to enhance the semantics for MIM. A line of studies [49, 38] using CLIP as a teacher to generate target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
2.0.2 Token aggregation methods.
@@
Token aggregation can conceptually be categorized as a token clustering method and usually aims for efficiency. Hard clustering methods like K-Means [30], K-Medoids [1], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) [23] enforce each data to belong to a single cluster exclusively. Bipartite matching [24] also aggregates data in a hard clustering manner, which optimizes pairs elements from two disjoint sets given objective function. Meanwhile, soft clustering is defined to let data belong to multiple clusters. LIT [33] employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods [37, 53, 40, 28] can be categorized into token aggregation methods; however, they intensely focused on compressing tokens to aim for a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We take the concept of token aggregation to address spatially noisy target tokens in token-level supervision, thereby enhancing the efficacy of MIM in terms of precision.
3 Pilot Study
In this section, we study the behavior of token representations generated by pre-trained models and their negative impacts on pre-training.
Spatial Inconsistency.
We begin by visualizing token representations from a vision-language model [35] to illustrate spatial inconsistency.
The inconsistent class predictions observed on each patch, as depicted in Fig. 1, reveal noisy token representations.
Impact of Spatial Inconsistency.
The spatially noisy token representations has the potential to lead to an underestimation of the capabilities of the vision-language models.
To understand the extent of these effects, we quantitatively assess various metrics at a dataset scale in Table 1. The quantitative analyses reveal that enhancing spatial coherence in the token representations results in improved performance, thus highlighting the importance of addressing spatial inconsistency.
Spatial Inconsistency in Representation Learning. The lack of spatial coherence in token representations have substantial negative impacts on the representation learning. Utilizing such noisy representations as targets can undermine the capabilities of the models. We employ representation learning to assess token aggregation methods to confirm the efficacy of handling spatial inconsistency via a refined context-aware method.
@@ 3 Pilot Study
| Token Aggregation | Zero-shot image classification () | Linear probing () | Averaged patch-wise cosine similarity with [CLS] |
| 26.5 | 73.2 | 0.53 | |
| ✓ | 30.8 (+4.3) | 77.6 (+3.2) | 0.56 (+0.3) |
3.1 Spatial Inconsistency
@@
We define the spatial inconsistency among token representations with an illustrative example. Given the input image in Fig. 1(a), we visualize the token-wise predictions without and with token aggregation in Fig. 1(b) and 1(c), respectively. Correct and incorrect tokens are marked in green and red, respectively, with a gradient to darker shades of red, indicating a more significant deviation from the true class. Despite the proximity and contextual similarity among tokens, wrong tokens in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (French Bulldog, Gorilla, Bull Mastiff, Hotdog, and Newfoundland Dog) while tokens in Fig. 1(c) show correct or relatively consistent predictions (french bulldog). Moreover, token aggregation for predictions yields improved accuracy, with 113 correct tokens out of 196 and a gap of 31 tokens compared to the counterpart. The comparison between the two prediction results highlights the spatial inconsistency among tokens from a pre-trained model, which could potentially disrupt representation learning when used as targets.
3.2 Impact of Spatial Inconsistency
To quantitatively assess the impact of spatial inconsistency at a dataset scale, we compute the token-wise ensemble prediction using a global pool, with and without token aggregations. Fundamentally, we predict class scores for each token and average these scores across all tokens within the given image. When predicting with token aggregation, we group semantically relevant tokens and average their representations group-wise prior to token-wise ensemble prediction. We employ CLIP-B/16 [35] for our study, so we aggregate 98 tokens for token aggregation, which occupy half of the entire tokens.
Table 1 reports the results of zero-shot image classification, linear probing, and averaged cosine similarity between the [CLS] on ImageNet-1K. Note that we do not perform any extra training here. Our results show that using aggregated tokens consistently exceeds the performances of those without token aggregation across all metrics. This suggests that aggregating contextually related tokens can address inconsistencies and lead to performance gains.
Note that the final metric we use is an alternative: a patch-wise continuous similarity metric. We measure the cosine similarity between the [CLS] token and each patch, which is also a patch-wise metric but continuous111 We further involve a continuous metric to assess spatial inconsistency, as it tracks network responses continuously compared with discrete classification metrics.. As observed in Table 1, the averaged similarities are computed to 0.56 vs. 0.53 for each case. This new metric, directly linked to the [CLS] token, suggests a more direct relationship between improved accuracy and reduced spatial inconsistency. The trend aligns with the other discrete metric, depicting patch-wise classification of 30.8 and 26.5 with and without token aggregation, respectively. The quantitative assessments, using both continuous and discontinuous metrics, demonstrate that token representations exhibit spatial inconsistency. Taking this further, we believe more effectively addressing inconsistency would yield a greater impact.
3.3 Spatial Inconsistency in Representation Learning
We now study the impact of token aggregation on representation learning. This analysis aims to highlight the spatial inconsistency in self-distillation signal could potentially diminish the quality of representations. Our comparison ranges distillations with the signals of 1) token-wise, 2) downsampling, 3) layer-wise bipartite matching during encoding, 4) superpixel, and 5) token morphing via bipartite matching. Note that token morphing will be introduced in §4.1, but we show its effectiveness in advance. We pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K for each configuration.
First, Table 2 exhibits distillation with token aggregation methods such as token morphing via bipartite matching, and superpixel clustering outperforms that of the token-wise distillation approach. Once again, this reveals that token representations indeed exhibit spatial inconsistency, as we claimed. Furthermore, token aggregation methods for contextually related tokens could manage this inconsistency. Since the downsampling approach aggregates tokens without context awareness, and the layer-wise bipartite matching approach risks harming intermediate representations during encoding, both approaches exhibit diminished returns. The findings highlight the significance of employing context-aware token aggregations that preserve token count while addressing spatial inconsistency.
Despite superpixel clustering’s promising results, our pilot study finds that our method learning with 1) superpixel clustering and 2) context-aware bipartite matching yields accuracies of 87.1% and 87.9%, respectively, over 100 epochs trained on ImageNet-100. Therefore, we do not employ a superpixel-based clustering for our method, which is inefficient as well.
4 Method - Dynamic Token Morphing
We observed that 1) pre-trained models often produce noisy, spatial inconsistency token-level targets for learning, closely linked with performance degradation; 2) token aggregation methods could handle spatial inconsistency but are insufficient as a supervisory signal; 3) a well-designed method is needed to consider context and reduce noise more effectively. Motivated by the observations, we introduce an advanced token aggregation method Dynamic Token Morphing (DTM) for masked image modeling and self-supervised pre-training. In DTM, we contextually aggregate tokens to derive random numbers of morphed tokens aiming to encapsulate diversified semantic information. The core idea of DTM is illustrated in Fig. 3, where the DTM module is straightforwardly added to the baseline MIM scheme. DTM is demonstrated for learning conventional token-level MIMs by aligning morphed tokens from online and target encoders.
4.1 Preliminary
Token encoding.
Given an image , we patchify the image into N patches .
We select positions of masked patches in a block-wise manner with a masking ratio so that .
We mask the image patches by replacing the image patches of the position in to a learnable mask token .
To this end, the input of the encoders becomes where for and for .
@@ 4.1 Preliminary
The masked patches are concatenated with a learnable CLS token and fed into the online encoder with a subsequent linear head and the target encoder , and become encoded online tokens and where and , respectively.
Token morphing. Differ to token aggregation methods, which prioritize efficiency via token reduction, token morphing specifically targets the spatial inconsistency in token representations by taking context into account. It connects contextually relevant tokens to smooth them without reducing the number of tokens (i.e., maintaining the token count). We define the process with the token morphing function based on the morphing schedule , a sequence of token numbers to morph. Here, is a general notation for a function that calculates similarity using a matching algorithm (e.g., bipartite matching or K-Means clustering) and returns a token morphing matrix , where means the number of token groups after morphing. Eventually, we can morph tokens following Eq. 3.
4.2 Overall Pipeline of Dynamic Token Morphing
We present a more advanced token morphing method, Dynamic Token Morphing (DTM), here. DTM’s dynamic nature stems from its simultaneous exploration of multiple and diverse token morphing cases. This design is based on the insight that morphing numerous tokens enhances the denoising effect while morphing fewer tokens retains detailed token representations. Essentially, the diversity in the process would yield significant benefits. Since morphing a fixed number of tokens may benefit only one side, DTM balances these advantages by morphing a diverse range of tokens. Furthermore, to achieve diversified morphed tokens, we divide the morphing process into multiple processes to create multiple morphed tokens. To this end, DTM encompasses three key components: 1) dynamic scheduler for token counts, 2) token morphing via scheduler, and 3) aligning morphed tokens. The overall framework of DTM is described in Fig. 4.
4.2.1 Dynamic scheduler.
DTM generates multiple morphed tokens to ensure diversity and an extensive range of token variations. To this end, we first sample the final number of morphed tokens to remain after token morphing and the iteration number from uniform distributions, where represents the minimum number of morphed tokens and denotes the maximum number of iteration for sampling. Then, we define a token count scheduler , a sequence of token numbers that determines the number of tokens to morph for each iteration dynamically. Rather than sampling a sequence of random numbers that satisfies , we simply divide by for constant counts:
4.2.2 Token morphing via dynamic scheduler.
The token morphing function is dynamically changed based on randomly sampled and . While our DTM is universal to any token matching methods, we employ bipartite matching [24] to search for contextually similar tokens efficiently. Specifically, we split tokens into two groups, with each token in the first group matched to its closest cosine similarity counterpart in the second group.